OpenStudy (anonymous):

In March, Isaiah and Freddie had the same amount of money in their savings accounts. In April, Isaiah deposited $210 into his account. Freddie increased the money in his account by 15%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in March? $_______

OpenStudy (anonymous):

x=Isaiah y= fred x+210=1.15(y) x+210=1.15x 210=.15x 1400=x

OpenStudy (anonymous):

yes, 210/0.15= 1400